OpenStudy (anonymous):

@JBrightman3 Which fraction falls between 5 -- 14 and 5 -- 7? ------------(Now the answers) 13 -- 14 ------------ 2 -- 7 ------------ 3 -- 14 ------------ 4 -- 7

OpenStudy (anonymous):

Well, 5/14 is about equal to 0.357, and 5/7 is about equal to 0.714. 13 / 14 = 0.929 2 / 7 = 0.286 3 / 14 = 0.214 4 / 7 = 0.571 So out of those numbers, which do you think is the correct answer?

OpenStudy (anonymous):

Think of it this way. \[\text{Which fraction falls between} \frac{5}{14} \text{and } \frac{10}{14}?\]\[\text{Choices } A.\frac{13}{14} \space B.\frac{4}{14} \space C.\frac{3}{14} \space D.\frac{8}{14}\]

OpenStudy (anonymous):

Uhh Mmmmm I think its D ?!?!

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

Yay!!! ^.^

OpenStudy (anonymous):

That's right! :)

OpenStudy (asnaseer):

Calcmathlete has shown the correct way of approaching this type of problem - make all the denominators the same - then just compare the numerators.

OpenStudy (anonymous):

Well, @asnaseer, you can find the answer either way. I just wanted to make it simpler for her. When you have the actual values of the fractions, it is easier to see the correct answer.

OpenStudy (anonymous):

Thanks guys!! ^.^

OpenStudy (anonymous):

Glad to help :)

OpenStudy (asnaseer):

@JBrightman3 you really think converting these fractions to /approx/ decimals is simpler?

OpenStudy (anonymous):

I just used a calculator and found each answer in less the a minute, so it was simple for me. :)

OpenStudy (anonymous):

@WhiteFlower, do you know how to change the denominators of fractions?

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

In the long run, it would be easier for you to solve the question that way, I just wasn't sure you've done those before. :)

OpenStudy (anonymous):

:) Yeah, thanks^^

OpenStudy (anonymous):

@JBrightman3 I do understand your point of view, but I think what he was trying to say was that that's the more appropriate way to do it because I've always learned it's better to do what can be done in your head without a calculator than with one. I remember very early on, I learned that kind of lesson that the mind can be faster than calculators and not to abuse calculators.

OpenStudy (anonymous):

I agree with you @Calcmathlete. I personally would have done it the way you did it, but I just didn't know if WhiteFlower knew how to yet. This one question is quite simple, so I didn't know if she knew how to or not.